# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ 2 Methodology
We assume that the entries of $\mathbf{P}$ are only partially observed, and furthermore, that the observed entries can be grouped into blocks. More specifically, suppose that we have $K$ sources and for any $i\in[K]$ we denote the index set of the entities contained in the $i$th source by $\mathcal{U}_{i}\subseteq[N]$. For ease of exposition we also require $\mathcal{U}_{i}\cap(\cup_{j\not=i}\mathcal{U}_{j})\not=\varnothing$ for all $i\in[K]$ as otherwise there exists some $i_{*}$ such that it is impossible to integrate observations from $\mathcal{U}_{i_{*}}$ with those from $\{\mathcal{U}_{j}\}_{j\not=i_{*}}$. We denote $n_{i}:=|\mathcal{U}_{i}|$ and the population matrix for the $i$th source by $\mathbf{P}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$. We then have
@@
We also allow for missing and corrupted observations in each source, i.e., for the $i$th source we only get to observe $\mathbf{P}^{(i)}_{s,t}+\mathbf{N}^{(i)}_{s,t}$ for all $\mathbf{\Omega}^{(i)}_{s,t}=1.$ Here $\mathbf{\Omega}^{(i)}\in\{0,1\}^{n_{i}\times n_{i}}$ indicates the indices of the observed entries and $\mathbf{N}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$ represent the random noise. In particular $\mathbf{\Omega}^{(i)}$ and $\mathbf{N}^{(i)}$ are both symmetric, and we assume the upper triangular entries of $\mathbf{\Omega}^{(i)}$ are i.i.d. Bernoulli random variables with success probability $q_{i}$ while the upper triangular entries of $\mathbf{N}^{(i)}$ are independent, mean-zero sub-Gaussian random variables with Orlicz-2 norm bounded by $\sigma_{i}=\max_{s,t\in[n_{i}]}\|\mathbf{N}^{(i)}_{s,t}\|_{\psi_{2}}$. For this model, the matrix
@@
is an unbiased estimate of $\mathbf{P}^{(i)}$, and thus a natural idea is to use the scaled leading eigenvectors $\hat{\mathbf{X}}^{(i)}=\hat{\mathbf{U}}^{(i)}(\hat{\bm{\Lambda}}^{(i)})^{1/2}$ as an estimate for $\mathbf{X}_{\mathcal{U}_{i}}$, where $\hat{\bm{\Lambda}}^{(i)}$ and $\hat{\mathbf{U}}^{(i)}$ contain the leading eigenvalues and the leading eigenvectors of $\mathbf{A}^{(i)}$, respectively. We now propose an algorithm to integrate and align $\{\hat{\mathbf{X}}^{(i)}\}_{i\in[K]}$ for recovery of the unobserved entries in $\mathbf{P}$.
@@ 2.1 Motivation of the algorithm
We first summarize the BONMI algorithm of Zhou et al. (2023). We start with the noiseless case for overlapping submatrices $\mathbf{P}^{(1)}$ and $\mathbf{P}^{(2)}$ to illustrate the main ideas. Our goal is to recover the unobserved entries in the white block of Figure 1; this is part of $\mathbf{P}_{\mathcal{U}_{1},\mathcal{U}_{2}}$.
Figure 1: A pair of overlapping observed submatrices
@@
and hence there exists $\mathbf{W}^{(1)},\mathbf{W}^{(2)}\in\mathcal{O}_{d}$ such that
@@
The same approach also extends to the case where the $\mathbf{P}^{(1)}$ and $\mathbf{P}^{(2)}$ are partially and noisily observed. More specifically, suppose we observe $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$ as defined in Eq. (2.1). We then obtain estimated latent positions $\hat{\mathbf{X}}^{(1)}$ for $\mathcal{U}_{1}$ and $\hat{\mathbf{X}}^{(2)}$ for $\mathcal{U}_{2}$ from $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$, respectively. To align $\hat{\mathbf{X}}^{(1)}$ and $\hat{\mathbf{X}}^{(2)}$, we solve the orthogonal Procrustes problem
$\mathbf{W}^{(1,2)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}^{(1)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\mathbf{O}-\hat{\mathbf{X}}^{(2)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\|_{F}$
and then estimate the unobserved block as part of $\hat{\mathbf{P}}_{\mathcal{U}_{1},\mathcal{U}_{2}}=\hat{\mathbf{X}}^{(1)}\mathbf{W}^{(1,2)}\hat{\mathbf{X}}^{(2)\top}.$
@@ 2.2 Chain-linked Multiple Matrix Integration (CMMI)
We now extend the ideas in Section 2.1 to a chain of overlapping submatrices. Suppose our goal is to recover the entries in the yellow block of Figure 2, given a collection $\{\mathbf{A}^{(i)}\}_{0\leq i\leq L}$ such that $\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\not=\varnothing$ for all $1\leq i\leq L$. Then for each pair $(\mathcal{U}_{i-1},\mathcal{U}_{i})$, we align the estimated latent position matrices $\hat{\mathbf{X}}^{(i-1)}$ and $\hat{\mathbf{X}}^{(i)}$ by solving the orthogonal Procrustes problem
$\mathbf{W}^{(i-1,i)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}^{(i-1)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}\mathbf{O}-\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}\|_{F}.$
Note that the solution of the orthogonal Procrustes problem between matrices $\hat{\mathbf{X}}^{(i-1)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}$ and $\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}$ is given by $\mathbf{W}_{1}\mathbf{W}_{2}^{\top}$ where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ contain the left and right singular vectors of $\hat{\mathbf{X}}^{(i-1)\top}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}\hat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}$, respecitvely; see Schönenmann (1966). $\hat{\mathbf{X}}^{(0)}$ and $\hat{\mathbf{X}}^{(L)}$ can be aligned by combining these $\{\mathbf{W}^{(i-1,i)}\}_{1\leq i\leq L}$ which then yields
$\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}=\hat{\mathbf{X}}^{(0)}\mathbf{W}^{(0,1)}\mathbf{W}^{(1,2)}\cdots\mathbf{W}^{(L-1,L)}\hat{\mathbf{X}}^{(L)\top}$
@@
Figure 2: A chain of overlapping observed submatrices
@@
    For $0\leq\ell\leq L$, obtain estimated latent position matrix for $\mathcal{U}_{i_{\ell}}$, denoted by $\hat{\mathbf{X}}^{(i_{\ell})}=\hat{\mathbf{U}}^{(i_{\ell})}(\hat{\bm{\Lambda}}^{(i_{\ell})})^{1/2}$, where $\hat{\mathbf{U}}^{(i_{\ell})}\in\mathbb{R}^{|\mathcal{U}_{i_{\ell}}|\times d}$ and the diagonal matrix $\hat{\bm{\Lambda}}^{(i_{\ell})}\in\mathbb{R}^{d\times d}$ contain the $d$ leading eigenvectors and eigenvalues of $\mathbf{A}^{(i_{\ell})}$, respectively.
@@
    $\mathbf{W}^{(i_{\ell-1},i_{\ell})}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{O}-\hat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\|_{F}.$
@@
    Compute $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}=\hat{\mathbf{X}}^{(i_{0})}\mathbf{W}^{(i_{0},i_{1})}\mathbf{W}^{(i_{1},i_{2})}\cdots\mathbf{W}^{(i_{L-1},i_{L})}\hat{\mathbf{X}}^{(i_{L})\top}$.
0:  $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}$.
@@
Compared to BONMI in Zhou et al. (2023), which handles only two submatrices, our proposed CMMI can combine all connected submatrices, where two submatrices $\mathbf{P}^{(i)}$ and $\mathbf{P}^{(j)}$ are said to be connected if there exists a path of overlapping submatrices between them. Indeed, for the example in Figure 2, BONMI can only recover the entries associated with pairs of overlapping submatrices, namely $\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{1}},\mathbf{P}_{\mathcal{U}_{1},\mathcal{U}_{2}},\dots,\mathbf{P}_{\mathcal{U}_{L-1},\mathcal{U}_{L}}$, while CMMI can recover the whole matrix $\mathbf{P}$. In general, BONMI only recovers $O(1/L)$ fraction of the entries recoverable by CMMI. Moreover, our theoretical results indicate that increasing $L$ has a minimal effect on the estimation error of CMMI (see Theorem 2), and simulations and real data experiments in Sections 4 and 5 show that accurate recovery is possible even when $L=20$. Our theoretical results also show that CMMI requires only minimal overlap between $\mathcal{U}_{i-1}$ and $\mathcal{U}_{i}$, e.g., $|\mathcal{U}_{i-1}\cap\mathcal{U}_{i}|$ can be as small as $d$, the embedding dimension of $\{\hat{\mathbf{X}}^{(i)}\}$; see Remark 6 and Section 4.3 for further discussion and experimental results.
@@
For more general cases encountered in practice, there may exist multiple chains to recover a given unobserved entry. We will present in Section A of the supplementary material several strategies to select and/or combine these chains. Although in certain cases, such as when considering a simple chain, CMMI is identical to the sequential integration approach SPSMC in Bishop and Yu (2014), CMMI offers a more optimal strategy in more complex scenarios by considering all overlapping pairs. In contrast, the restriction to sequential integration imposes limitations on SPSMC, and how to determine a valid and effective sequential integration order is an unresolved issue in Bishop and Yu (2014). Furthermore, our theoretical results are significantly stronger than those in Bishop and Yu (2014) and Zhou et al. (2023); see Section 3.1 for further discussion.

### v4

@@ 2 Methodology
We assume that the entries of $\mathbf{P}$ are only partially observed, and furthermore, that the observed entries can be grouped into blocks. More specifically, suppose that we have $K$ sources and for any $i\in[K]$ we denote the index set of the entities contained in the $i$th source by $\mathcal{U}_{i}\subseteq[N]$. For ease of exposition we also require $\mathcal{U}_{i}\cap(\cup_{j\not=i}\mathcal{U}_{j})\not$$\varnothing$ for all $i\in[K]$ as otherwise there exists some $i_{*}$ such that it is impossible to integrate observations from $\mathcal{U}_{i_{*}}$ with those from $\{\mathcal{U}_{j}\}_{j\not=i_{*}}$. We denote $n_{i}:=|\mathcal{U}_{i}|$ and the population matrix for the $i$th source by $\mathbf{P}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$. We then have
@@
We also allow for missing and corrupted observations in each source, i.e., for the $i$th source we only get to observe $\mathbf{P}^{(i)}_{s,t}+\mathbf{N}^{(i)}_{s,t}$ for all $\mathbf{\Omega}^{(i)}_{s,t}=1.$ Here $\mathbf{\Omega}^{(i)}\in\{0,1\}^{n_{i}\times n_{i}}$ indicates the indices of the observed entries and $\mathbf{N}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$ represents the random noise. In particular $\mathbf{\Omega}^{(i)}$ and $\mathbf{N}^{(i)}$ are both symmetric, and we assume the upper triangular entries of $\mathbf{\Omega}^{(i)}$ are i.i.d. Bernoulli random variables with success probability $q_{i}$ while the upper triangular entries of $\mathbf{N}^{(i)}$ are independent, mean-zero sub-Gaussian random variables with Orlicz-2 norm bounded by $\sigma_{i}:=\max_{s,t\in[n_{i}]}\|\mathbf{N}^{(i)}_{s,t}\|_{\psi_{2}}$. For this model, the matrix
@@
is an unbiased estimate of $\mathbf{P}^{(i)}$, and thus a natural idea is to use the scaled leading eigenvectors $\widehat{\mathbf{X}}^{(i)}=\widehat{\mathbf{U}}^{(i)}(\widehat{\bm{\Lambda}}^{(i)})^{1/2}$ as an estimate for $\mathbf{X}_{\mathcal{U}_{i}}$, where $\widehat{\bm{\Lambda}}^{(i)}$ and $\widehat{\mathbf{U}}^{(i)}$ contain the leading eigenvalues and the leading eigenvectors of $\mathbf{A}^{(i)}$, respectively. Note that in practice we use the empirical observed proportion $\widehat{q}_{i}$ in place of $q_{i}$ when constructing $\mathbf{A}^{(i)}$. In particular, our evaluation of the algorithm’s performance in the numerical experiments is based entirely on $\widehat{q}_{i}$. In contrast we assume that $q_{i}$ is known in our theoretical analysis. This is done, both for ease of exposition and without loss of generality, as $|\widehat{q}_{i}-q_{i}|=O_{p}(n_{i}^{-1})$ for all $i$ and thus has no impact on the theoretical results stated in Section 3. We now propose an algorithm to integrate and align $\{\widehat{\mathbf{X}}^{(i)}\}_{i\in[K]}$ for recovery of the unobserved entries in $\mathbf{P}$.
@@ 2.1 Motivation of the algorithm
We first summarize the BONMI algorithm of Zhou et al. (2023). We start with the noiseless case for two overlapping submatrices $\mathbf{P}^{(1)}$ and $\mathbf{P}^{(2)}$ to illustrate the main ideas. Our goal is to recover the unobserved entries for the white block in Figure 1; this is part of $\mathbf{P}_{\mathcal{U}_{1},\mathcal{U}_{2}}$.
Figure 1: a pair of overlapping observed submatrices.
@@
and hence there exist $\mathbf{W}^{(1)},\mathbf{W}^{(2)}\in\mathcal{O}_{d}$ such that
@@
The same approach also extends to the case where the $\mathbf{P}^{(1)}$ and $\mathbf{P}^{(2)}$ are partially and noisily observed. More specifically, suppose we observe $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$ as defined in Eq. (2.1). We then obtain estimated latent positions $\widehat{\mathbf{X}}^{(1)}$ for $\mathcal{U}_{1}$ and $\widehat{\mathbf{X}}^{(2)}$ for $\mathcal{U}_{2}$ from $\mathbf{A}^{(1)}$ and $\mathbf{A}^{(2)}$, respectively. To align $\widehat{\mathbf{X}}^{(1)}$ and $\widehat{\mathbf{X}}^{(2)}$, we solve the orthogonal Procrustes problem
$\mathbf{W}^{(1,2)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(1)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(2)}_{\langle\mathcal{U}_{1}\cap\mathcal{U}_{2}\rangle}\|_{F}$
and then estimate the unobserved block as part of $\widehat{\mathbf{P}}_{\mathcal{U}_{1},\mathcal{U}_{2}}=\widehat{\mathbf{X}}^{(1)}\mathbf{W}^{(1,2)}\widehat{\mathbf{X}}^{(2)\top}.$
@@ 2.2 Chain-linked Multiple Matrix Integration (CMMI)
We now extend the ideas in Section 2.1 to a chain of overlapping submatrices. Suppose our goal is to recover the entries in the yellow block in Figure 2. Given a collection $\{\mathbf{A}^{(i)}\}_{0\leq i\leq L}$ such that $\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\not$$\varnothing$ for all $1\leq i\leq L$. Then for each pair $(\mathcal{U}_{i-1},\mathcal{U}_{i})$, we align the estimated latent position matrices $\widehat{\mathbf{X}}^{(i-1)}$ and $\widehat{\mathbf{X}}^{(i)}$ by solving the orthogonal Procrustes problem
$\mathbf{W}^{(i-1,i)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(i-1)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}\|_{F}.$
Note that the solution of the orthogonal Procrustes problem between matrices $\widehat{\mathbf{X}}^{(i-1)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}$ and $\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}$ is given by $\mathbf{M}_{1}\mathbf{M}_{2}^{\top}$ where $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ contain the left and right singular vectors of $\widehat{\mathbf{X}}^{(i-1)\top}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i-1}\cap\mathcal{U}_{i}\rangle}$, respectively (Schönenmann, 1966). $\widehat{\mathbf{X}}^{(0)}$ and $\widehat{\mathbf{X}}^{(L)}$ can be aligned by combining these $\{\mathbf{W}^{(i-1,i)}\}_{1\leq i\leq L}$, which then yields
$\widehat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}=\widehat{\mathbf{X}}^{(0)}\mathbf{W}^{(0,1)}\mathbf{W}^{(1,2)}\cdots\mathbf{W}^{(L-1,L)}\widehat{\mathbf{X}}^{(L)\top}$
@@
Figure 2: a chain of overlapping observed submatrices.
@@
    For $0\leq\ell\leq L$, obtain estimated latent position matrix for $\mathcal{U}_{i_{\ell}}$, denoted by $\widehat{\mathbf{X}}^{(i_{\ell})}=\widehat{\mathbf{U}}^{(i_{\ell})}(\widehat{\bm{\Lambda}}^{(i_{\ell})})^{1/2}$, where $\widehat{\mathbf{U}}^{(i_{\ell})}\in\mathbb{R}^{|\mathcal{U}_{i_{\ell}}|\times d}$ and the diagonal matrix $\widehat{\bm{\Lambda}}^{(i_{\ell})}\in\mathbb{R}^{d\times d}$ contain the $d$ leading eigenvectors and eigenvalues of $\mathbf{A}^{(i_{\ell})}$, respectively.
@@
    $\mathbf{W}^{(i_{\ell-1},i_{\ell})}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\|_{F}.$
@@
    Compute $\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}=\widehat{\mathbf{X}}^{(i_{0})}\mathbf{W}^{(i_{0},i_{1})}\mathbf{W}^{(i_{1},i_{2})}\cdots\mathbf{W}^{(i_{L-1},i_{L})}\widehat{\mathbf{X}}^{(i_{L})\top}$.
[…]
For choosing $d$ in Algorithm 1, in practice we can first examine the eigenvalues of $\{\mathbf{A}^{(i)}\}$ to select the individual $\{d_{i}\}$, and then set $d=\max_{i}d_{i}$ to ensure that it captures all relevant signal components. One widely-used approach for selecting an individual $d_{i}$ is to inspect the scree plot for $\mathbf{A}^{(i)}$ and identify a “elbow” separating the signal eigenvalues from the noise eigenvalues. An example of this approach is the automated dimensionality selection procedure in Zhu and Ghodsi (2006) which maximizes a profile likelihood function. Other examples include residual subsampling (Han et al., 2023), eigenvalue ratio tests (Ahn and Horenstein, 2013), and inference based on empirical eigenvalue distributions (Onatski, 2010).
@@
Compared to BONMI in Zhou et al. (2023), which handles only two overlapping submatrices at a time, our proposed CMMI can actually combine all connected submatrices, where two submatrices $\mathbf{P}^{(i)}$ and $\mathbf{P}^{(j)}$ are said to be connected if there exists a path of overlapping submatrices between them. Indeed, for the example in Figure 2, BONMI can only recover the entries associated with pairs of overlapping submatrices, namely $\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{1}},\mathbf{P}_{\mathcal{U}_{1},\mathcal{U}_{2}},\dots,\mathbf{P}_{\mathcal{U}_{L-1},\mathcal{U}_{L}}$, while CMMI can recover the whole matrix $\mathbf{P}$. In general, BONMI only recovers $O(1/L)$ fraction of the entries recoverable by CMMI. Moreover, our theoretical results indicate that increasing $L$ has a minimal effect on the estimation error of CMMI (see Theorem 2), and simulations and real data experiments in Sections 4 and 5 show that accurate recovery is possible even when $L=20$. Our theoretical results also show that CMMI requires only minimal overlap between $\mathcal{U}_{i-1}$ and $\mathcal{U}_{i}$, e.g., $|\mathcal{U}_{i-1}\cap\mathcal{U}_{i}|$ can be as small as $d$, the embedding dimension of $\{\widehat{\mathbf{X}}^{(i)}\}$; see Remark 5 for further discussion, and Section A.2 of the supplementary material for corresponding experimental results.
@@
For more general cases encountered in practice, the structures of the observed submatrices can be more complex than simple chains. In Section B of the supplementary material, we extend the basic CMMI algorithm from Algorithm 1 to integrate connected submatrices with arbitrarily complex structures. Although in certain cases, such as a simple chain, CMMI is identical to the sequential integration approach SPSMC in Bishop and Yu (2014), CMMI offers a more refined strategy in more complex scenarios by considering all overlapping pairs, as the restriction to sequential integration imposes limitations on SPSMC, and how to determine an effective integration order is unresolved in Bishop and Yu (2014).
This idea of first obtaining individual estimates and then sequentially aligning them to obtain a global estimate also appears in the Spectral-Stitching algorithm in Chen et al. (2016) where the goal is to determine the community membership of each vertex in a graph with two communities. More specifically, the Spectral-Stiching algorithm first partitions the vertex set into several overlapping subsets of size $n$ such that any two adjacent subsets share $n/2$ common vertices. It then applies spectral methods separately to each subgraph to obtain community estimates. Since the community labels obtained from different subgraphs may be inconsistent, the algorithm sequentially stitches the individual community estimates together using majority voting.